# Electronic modes induced by spin and charge perturbations in Mott and Kondo insulators

### v1

@@ A.3 Spectral function of nonzero-$S$ states
The electron-addition and electron-removal spectral functions of a state of $S=n$ and $S^{z}=m$ for $H=0$ can be expressed as
@@
with $(T^{\frac{1}{2}}_{\frac{1}{2}},T^{\frac{1}{2}}_{-\frac{1}{2}})=(-c_{{\bm{k}},\downarrow},c_{{\bm{k}},\uparrow})$ and $Z_{n}=Z_{n,m}$, where $i\in(n\pm\frac{1}{2})$ means that $|i\rangle$ has $S=n\pm\frac{1}{2}$, and the quantities with subscript $i$ indicate those for $|i\rangle$; the excitation energy of $|i\rangle$ from $|n,m\rangle$ is denoted as $\epsilon_{i}$. Here, $\epsilon_{i}$ does not depend on $S^{z}$ of $|i\rangle$ in a spin-SU(2) multiplet, and $\langle n^{\prime},m^{\prime}|T^{\frac{1}{2}}_{s^{z}}|n,m\rangle$ can be nonzero only for $n^{\prime}=n\pm\frac{1}{2}$ and $m^{\prime}=m+s^{z}$.
@@ A.4 Relation between spectral functions before and after Shiba transformation
The spectral function of a state $|X\rangle$ for $\lambda(=c,f)$ electrons of spin $\sigma(=\uparrow,\downarrow)$ is defined [Eq. (10)] as
@@
for ${\cal H}(-U)$, where $|X^{\prime}\rangle$ is the state of $\eta=j$ and $\eta^{z}=m$. Here, the excitation energy of $|i\rangle$ from $|X\rangle$ or $|X^{\prime}\rangle$ is denoted as $\epsilon_{i}$, and ${\cal H}(U)$ and ${\cal H}(-U)$ denote the Hamiltonians before and after the Shiba transformation, respectively [Eqs. (100) and (V.5)]. To distinguish the spectral functions before and after the transformation, the spectral function after the Shiba transformation is denoted as ${\tilde{A}}^{\lambda,\sigma}_{X^{\prime}}({\bm{k}},\omega)$. The $\uparrow$-spin electron spectral function is transformed as
@@
Thus, the spectral functions of a state of $S=j$, $S^{z}=m$, and $\eta=0$ at half filling ($\eta^{z}=0$) at $\mu=0$ for ${\cal H}(U)$ can be expressed using those of the states of $\eta=j$, $\eta^{z}=\pm m$, and $S=0$ at $H=0$ for ${\cal H}(-U)$ as Eqs. (230)–(A.4); in particular, the spectral functions of the magnetized ground state at half filling for ${\cal H}(U)$ can be expressed using those of the doped ground states for ${\cal H}(-U)$.
@@ B.1 Effective states and energies in systems of weakly coupled unit cells
$N_{\rm e}$ $S$ $S^{z}$ $\eta$ $\eta^{z}$ $k_{\perp}$ Eigenstate Energy
@@
0 0 0 1 -1 0 $|{\rm V}\rangle=|0,0\rangle$ ${\cal E}_{\rm V}=\frac{U}{2}$
@@
  • The state with $|\alpha_{1}\rangle$ and $|\alpha_{2}\rangle$ at sites 1 and 2, respectively, in a dimer is denoted as $|\alpha_{1},\alpha_{2}\rangle$.
@@
  • The state with $|\alpha_{c}\rangle$ in the conduction orbital and $|\alpha_{f}\rangle$ in the localized orbital is denoted as $|\alpha_{c},\alpha_{f}\rangle$.
@@
  • The state with $|\alpha_{c}\rangle$ in the conduction orbital and $|\alpha_{f}\rangle$ in the localized orbital is denoted as $|\alpha_{c},\alpha_{f}\rangle$.
@@
$|{\rm GS}\rangle=\prod_{i=1}^{N_{\rm u}}|{\cal G}\rangle_{i},$ (236)
$|X({\bm{k}}_{\parallel})\rangle=\frac{1}{\sqrt{N_{\rm u}}}\sum_{i=1}^{N_{\rm u}}e^{i{\bm{k}}_{\parallel}\cdot{\bm{r}}_{i}}|X\rangle_{i}\prod_{j(\neq i)}^{N_{\rm u}}|{\cal G}\rangle_{j},$ (237)
@@ B.2 Symmetry of effective parameters
In the HM, the energies of the bonding electronic state with $k_{\perp}=0$ and antibonding electronic state with $k_{\perp}=\pi$ are interchanged by the transformation of $t_{\perp}\rightarrow-t_{\perp}$;
@@
In the HM and symmetric PAM at $\mu=0$, $\hat{\eta}$$^{\pm}$ shifts a momentum $\bm{k}$} by $\bm{\pi}$} because of $(-)^{i}$ [Eqs. (18) and (22)];

### v2

@@ A.3 Spectral function of nonzero-$S$ states
The electron-addition and electron-removal spectral functions of an eigenstate of the Hamiltonian with $S=n$ and $S^{z}=m$ for $H=0$ can be expressed as
@@
with $(T^{\frac{1}{2}}_{\frac{1}{2}},T^{\frac{1}{2}}_{-\frac{1}{2}})=(-c_{{\bm{k}},\downarrow},c_{{\bm{k}},\uparrow})$ and $Z_{n}=Z_{n,m}$, where $i\in(n\pm\frac{1}{2})$ means that $|i\rangle$ has $S=n\pm\frac{1}{2}$, and the quantities with subscript $i$ indicate those for $|i\rangle$; the excitation energy of $|i\rangle$ from $|n,m\rangle$ is denoted by $\epsilon_{i}$. Here, $\epsilon_{i}$ does not depend on $S^{z}$ of $|i\rangle$ in a spin-SU(2) multiplet, and $\langle n^{\prime},m^{\prime}|T^{\frac{1}{2}}_{s^{z}}|n,m\rangle$ can be nonzero only for $n^{\prime}=n\pm\frac{1}{2}$ and $m^{\prime}=m+s^{z}$.
@@ A.4 Relation between spectral functions before and after Shiba transformation
The spectral function of an eigenstate of the Hamiltonian $|X\rangle$ for $\lambda(=c,f)$ electrons of spin $\sigma(=\uparrow,\downarrow)$ can be expressed [Eq. (12)] as
@@
for ${\cal H}(-U)$, where $|X^{\prime}\rangle$ is the state of $\eta=j$ and $\eta^{z}=m$. Here, the excitation energy of $|i\rangle$ from $|X\rangle$ or $|X^{\prime}\rangle$ is denoted by $\epsilon_{i}$, and ${\cal H}(U)$ and ${\cal H}(-U)$ denote the Hamiltonians before and after the Shiba transformation, respectively [Eqs. (102) and (V.5)]. To distinguish the spectral functions before and after the transformation, the spectral function after the Shiba transformation is denoted by ${\tilde{A}}^{\lambda,\sigma}_{X^{\prime}}({\bm{k}},\omega)$. The $\uparrow$-spin electron spectral function is transformed as
@@
Thus, the spectral functions of an eigenstate with $S=j$, $S^{z}=m$, and $\eta=0$ at half filling ($\eta^{z}=0$) at $\mu=0$ for ${\cal H}(U)$ can be expressed using those of the eigenstates with $\eta=j$, $\eta^{z}=\pm m$, and $S=0$ at $H=0$ for ${\cal H}(-U)$ as Eqs. (232)–(A.4); in particular, the spectral functions of the magnetized ground state at half filling for ${\cal H}(U)$ can be expressed using those of the doped ground states for ${\cal H}(-U)$.
@@ B.1 Effective states and energies in systems of weakly coupled unit cells
$N_{\rm e}$ $S$ $S^{z}$ $\eta$ $\eta^{z}$ $k_{\perp}$ Eigenstates Energies
@@
0 0 0 1 1-1 0 $|{\rm V}\rangle=|0,0\rangle$ ${\cal E}_{\rm V}=\frac{U}{2}$
@@
  • The state with $|\alpha_{1}\rangle$ and $|\alpha_{2}\rangle$ at sites 1 and 2, respectively, in a dimer is denoted by $|\alpha_{1},\alpha_{2}\rangle$.
@@
  • The state with $|\alpha_{c}\rangle$ in the conduction orbital and $|\alpha_{f}\rangle$ in the localized orbital is denoted by $|\alpha_{c},\alpha_{f}\rangle$.
@@
  • The state with $|\alpha_{c}\rangle$ in the conduction orbital and $|\alpha_{f}\rangle$ in the localized orbital is denoted by $|\alpha_{c},\alpha_{f}\rangle$.
@@
$|{\rm GS}\rangle=\prod_{i}|{\cal G}\rangle_{i},$ (238)
$|X\rangle_{{\bm{k}}_{\parallel}}=\frac{1}{\sqrt{N_{\rm u}}}\sum_{i}e^{i{\bm{k}}_{\parallel}\cdot{\bm{r}}_{i}}|X\rangle_{i}\prod_{j(\neq i)}|{\cal G}\rangle_{j},$ (239)
@@ B.2 Symmetry of effective parameters
In the HM, the energies of the bonding electronic state with $k_{\perp}=0$ and antibonding electronic state with $k_{\perp}=\pi$ are interchanged by the transformation of $t_{\perp}\rightarrow-t_{\perp}$,
@@
In the HM and symmetric PAM at $\mu=0$, $\hat{\eta}$$^{\pm}$ shifts a momentum $\bm{k}$} by $\bm{\pi}$} because of $(-)^{i}$ [Eqs. (20) and (24)],